### v3

@@ C.2 Data Granularity Justification
C.3 Automated Data Annotation
@@ Appendix D Annotation Data Quality Control
D.1 Validation Data Filtering
@@ D.2 Annotator Demography
D.3 Validation Guideline
D.3.1 Task Definition
@@ D.3.2 Validation Procedure
Table 10: Entity type definitions used for validation.
@@ D.4 Agreement Computation
$\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}.$ (7)
Following landis1977measurement, our $\kappa=0.81$ indicates substantial agreement.
@@ Appendix E Evaluation Metrics
To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work sharma2023financial. Macro-level evaluation computes precision, recall, and F1 scores independently for each BIO-concept label derived from the US-GAAP taxonomy, and then averages them without weighting. This ensures that each concept, including rare or infrequent ones, contributes equally to the final score, making it especially suitable for domains with skewed label distributions. In contrast, micro-level evaluation aggregates token-level true positives, false positives, and false negatives across all labels before computing precision, recall, and F1. This approach emphasizes the model’s overall tagging accuracy by treating every token equally and thus better reflects performance on frequent concepts. Together, these two metrics provide a balanced view of both per-concept performance and overall tagging quality.
@@
$\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$ (8)
$\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$ (9)
$\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (10)
@@
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$ (11)
@@ Appendix F Evaluation Models Details
  • Closed-source LLMs: We include GPT-4o hurst2024gpt, accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks. Although model size details are undisclosed, GPT-4o serves as an upper-bound reference in our benchmark.
  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) liu2024deepseek, DeepSeek-R1-Distill-Qwen (32B) deepseekai2025deepseekr1incentivizingreasoningcapability, and multiple variants of Qwen3  qwen3technicalreport(ranging from 0.6B to 32B). We also include Llama-3.3, Llama-3.2, and 3.1 variants  grattafiori2024llama (70B, 3B, and 8B), Llama-4-Scout-17B-16E-Instruct meta2025llama(109B), as well as Google’s gemma-2-27B  gemma_2024, to ensure architectural diversity and scalability comparison. These models are primarily instruction-tuned and optimized for general-purpose NLP tasks.
  • Financial-specific LLMs: We evaluate Fino1-8B  qian2025fino1, a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures. This category allows us to assess the benefits of domain adaptation in complex tagging and reasoning tasks.
  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  DBLP:journals/corr/abs-1810-04805, FinBERT  araci2019finbert, and SECBERT  loukas2022finer. These models have been widely used in prior financial NLP tasks and allow for a comparative analysis between generative LLMs and traditional pretrained models in terms of domain understanding and structured output capability.
@@
Table 11: Model categories and corresponding repositories.
@@ Appendix G The details for the fine-tuning PTMs
G.1 Training data collection and processing
@@
Table 12: Financial report statistics summary for raw training data
@@
Table 13: Statistics of training data (tokens calculated with “cl100k_base” tokenizer, $\pm$ standard deviation).
@@
However, to align with the extreme classification format used in previous XBRL tagging benchmarks, we directly adopt the US-GAAP tags as entity labels, annotating each token in sentences and tables using the BIO scheme. Specifically, B denotes the beginning of an entity phrase, I marks the continuation (inside) of an entity phrase, and O indicates tokens outside of any entity. As shown in Figure 8, “4.9” and “4.5” are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”). To comprehensively cover all US-GAAP tags, we combine the entire set of 17,688 tags from the US-GAAP 2023 and 2024 taxonomy with the BIO labeling scheme to construct an extreme classification label space, resulting in 34,777 unique entity labels ($2\times 17388+1$).
Figure 8: An example for training set annotation with BIO scheme.
@@
G.2 Training settings
We fine-tune three pretrained models, BERT-large DBLP:journals/corr/abs-1810-04805, FinBERT araci2019finbert, and SECBERT loukas2022finer, on our training set using the HuggingFace Transformers library. All models are trained with a batch size of 4, a learning rate of 3e-5, and for 20 epochs. Optimization is performed using AdamW without gradient accumulation or early stopping. Token classification heads are randomly initialized and trained jointly with the base encoder. Input sequences are tokenized with a maximum length of 512, and labels are aligned at the sub-token level following the BIO tagging scheme. Loss is computed only on the first sub-token of each word to avoid misalignment bias.
@@ Appendix H Retrieval Results in FinCL subtask
Table 14: Acc@k retrieval performance on the FinCL task. FWC: fixed window-based; SAC: structure-aware.
@@ Appendix I Error cases
Table 15: Error cases from DeepSeek-V3 rerank.
@@
Figure 9: Prompt template used for the FinNI subtask.
@@
Figure 10: Prompt template used for the Reranking stage in the FinCL subtask.
@@
Figure 11: Prompt template used for ablation study.

### v4

@@ C.2. Data Granularity Justification
C.3. Automated Data Annotation
@@ Appendix D Annotation Data Quality Control
D.1. Validation Data Filtering
@@ D.2. Annotator Demography
D.3. Validation Guideline
D.3.1. Task Definition
@@ D.3.2. Validation Procedure
Table 9. Entity type definitions used for validation.
@@ D.4. Agreement Computation
(6) $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}.$
Following Landis and Koch (1977), our $\kappa=0.81$ indicates substantial agreement.
@@ Appendix E Evaluation Metrics
To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work (Sharma et al., 2023). Macro-level evaluation computes precision, recall, and F1 scores independently for each BIO-concept label derived from the US-GAAP taxonomy, and then averages them without weighting. This ensures that each concept, including rare or infrequent ones, contributes equally to the final score, making it especially suitable for domains with skewed label distributions. In contrast, micro-level evaluation aggregates token-level true positives, false positives, and false negatives across all labels before computing precision, recall, and F1. This approach emphasizes the model’s overall tagging accuracy by treating every token equally and thus better reflects performance on frequent concepts. Together, these two metrics provide a balanced view of both per-concept performance and overall tagging quality.
@@
(7) $\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$
(8) $\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$
(9) $\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
@@
(10) $\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$
@@ Appendix F Evaluation Models Details
  • Closed-source LLMs: We include GPT-4o (Hurst et al., 2024), accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks. Although model size details are undisclosed, GPT-4o serves as an upper-bound reference in our benchmark.
  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) (Liu et al., 2024), DeepSeek-R1-Distill-Qwen (32B) (DeepSeek-AI, 2025), and multiple variants of Qwen3  (Team, 2025)(ranging from 0.6B to 32B). We also include Llama-3.3, Llama-3.2, and 3.1 variants  (Grattafiori et al., 2024) (70B, 3B, and 8B), Llama-4-Scout-17B-16E-Instruct (Meta, 2025)(109B), as well as Google’s gemma-2-27B  (Team, 2024), to ensure architectural diversity and scalability comparison. These models are primarily instruction-tuned and optimized for general-purpose NLP tasks.
  • Financial-specific LLMs: We evaluate Fino1-8B  (Qian et al., 2025), a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures. This category allows us to assess the benefits of domain adaptation in complex tagging and reasoning tasks.
  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  (Devlin et al., 2018), FinBERT  (Araci, 2019), and SECBERT  (Loukas et al., 2022). These models have been widely used in prior financial NLP tasks and allow for a comparative analysis between generative LLMs and traditional pretrained models in terms of domain understanding and structured output capability.
@@
Table 10. Model categories and corresponding repositories.
@@ Appendix G The details for the fine-tuning PTMs
G.1. Training data collection and processing
@@
Table 11. Financial report statistics summary for raw training data
@@
Table 12. Statistics of training data (tokens calculated with “cl100k_base” tokenizer, $\pm$ standard deviation).
@@
However, to align with the extreme classification format used in previous XBRL tagging benchmarks, we directly adopt the US-GAAP tags as entity labels, annotating each token in sentences and tables using the BIO scheme. Specifically, B denotes the beginning of an entity phrase, I marks the continuation (inside) of an entity phrase, and O indicates tokens outside of any entity. As shown in Figure 8, ”4.9” and ”4.5” are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”). To comprehensively cover all US-GAAP tags, we combine the entire set of 17,688 tags from the US-GAAP 2023 and 2024 taxonomy with the BIO labeling scheme to construct an extreme classification label space, resulting in 34,777 unique entity labels ($2\times 17388+1$).
Figure 8. An example for training set annotation with BIO scheme.
@@
G.2. Training settings
We fine-tune three pretrained models, BERT-large (Devlin et al., 2018), FinBERT (Araci, 2019), and SECBERT (Loukas et al., 2022), on our training set using the HuggingFace Transformers library. All models are trained with a batch size of 4, a learning rate of 3e-5, and for 20 epochs. Optimization is performed using AdamW without gradient accumulation or early stopping. Token classification heads are randomly initialized and trained jointly with the base encoder. Input sequences are tokenized with a maximum length of 512, and labels are aligned at the sub-token level following the BIO tagging scheme. Loss is computed only on the first sub-token of each word to avoid misalignment bias.
@@ Appendix H Retrieval Results in FinCL subtask
Table 13. Acc@k retrieval performance on the FinCL task. FWC: fixed window-based; SAC: structure-aware.
@@ Appendix I Error cases
Table 14. Error cases from DeepSeek-V3 rerank.
@@
Figure 9. Prompt template used for the FinNI subtask.
@@
Figure 10. Prompt template used for the Reranking stage in the FinCL subtask.
@@
Figure 11. Prompt template used for ablation study.